# Trustless Autonomy: Understanding Motivations, Benefits and Governance Dilemma in Self-Sovereign Decentralized AI Agents

### v1

@@ Abstract.
The recent trend of self-sovereign Decentralized AI Agents (DeAgents) combines Large Language Model (LLM)-based AI agents with decentralization technologies such as blockchain smart contracts and trusted execution environments (TEEs). These tamper-resistant trustless substrates allow agents to achieve self-sovereignty through ownership of cryptowallet private keys and control of digital assets and social media accounts. DeAgent eliminates centralized control and reduces human intervention, addressing key trust concerns inherent in centralized AI systems. However, given ongoing challenges in LLM reliability such as hallucinations, this creates paradoxical tension between trustlessness and unreliable autonomy. This study addresses this empirical research gap through interviews with DeAgents stakeholders—experts, founders, and developers—to examine their motivations, benefits, and governance dilemmas. The findings will guide future DeAgents system and protocol design and inform discussions about governance in sociotechnical AI systems in the future agentic web.
@@ 1. Introduction
Deploying DeAgents on trustless secure enclaves, such as smart contracts and TEEs, inherently creates paradoxical tensions around governance risks due to diminished human oversight. Proponents envision that self-sovereign DeAgents can help moderate decentralized autonomous organizations (DAO) (Hinkes, 2024), manage financial portfolios (Ante, 2024), and etc., aiming to mitigate human error and centralized corruption (Trequattrini et al., 2024; Odilla, 2023). For example, ai16z is the first AI venture capital fund that is fully managed by DeAgent with recommendations from DAO members (ai1, 2024). However, while trustless infrastructure can guarantee tamper-resistant autonomous execution (Jauernig et al., 2020), the AI agent itself may not be inherently trustworthy (Liu et al., 2024). Large language models are prone to errors (Yadkori et al., 2024), biases (Rettenberger et al., 2025), deceptions (Hubinger et al., 2024), hallucinations (Huang et al., 2025; Xu et al., 2025a), and misalignment (Shen et al., 2023). The additional challenge of ”locking” AI agents in secure enclaves to prevent external tampering arguably amplifies potential harms if the agents’ decisions are flawed or hacked (Patlan et al., 2025)—especially without a kill switch (Seneviratne, 2024). This tension between trustlessness and autonomy fundamentally challenges the governability (Hu et al., 2025), accountability (Novelli et al., 2024) and responsibility (Dignum, 2019) of sociotechnical AI systems (Kudina and van de Poel, 2024; Dobbe and Wolters, 2024; Feng et al., 2025a).
@@
  • RQ2 (Challenges and Governance Implications): What challenges arise in the real-world deployment of DeAgents, and how do stakeholders navigate governance, oversight, and accountability within a trustless autonomous system?
@@
  • We presented empirical findings showing that the push for self-sovereign AI agents on trustless substrates is driven by sociotechnical / political factors (Fratini, 2025) like decentralization, privacy concerns, censorship resistance, and community ownership—even as AI reliability issues remain unresolved.
  • We identified the paradoxical governance dilemma that arises when potentially unreliable AI agents operate within trustless infrastructure, revealing the complex tensions between trusting humans versus trusting AI in delegating agency to autonomous systems.
  • We proposed design and governance implications that emphasize thoughtful protocol design for agentic web of trust (Chaffer, 2025; Habler et al., 2025), with flexible fail-safes serving as essential bridges between trustlessness and genuine trustworthiness.
@@ 2.1. Terminology used in this paper
The capacity of an agent to perceive, decide, and act toward its goals without continuous human oversight, also known as ”self-governance.” However, higher powers may still exist that can assume control over the agent (Hui, 2024).
@@ 2.2. LLM-based Agents
Recent advancements in Large Language Models have given rise to autonomous agents (Wang et al., 2024; Xi et al., 2023) capable of performing various real-world tasks (Xu et al., 2024), such as automatic social media marketing (Deshmukh et al., 2025), scientific discovery (Ren et al., 2025; Gridach et al., 2025) and financial trading (Ding et al., 2024). These LLM-based agents typically share a common architectural framework (Luo et al., 2025) that includes profile definition, memory mechanisms (Zhang et al., 2024), planning capabilities (Huang et al., 2024; He et al., 2025), action execution via context protocol and external tools (Hou et al., 2025), inter-agent collaboration (Mou et al., 2024), and agent evolution (Dai et al., 2024). Open-source agentic frameworks such as LangChain and AutoGen (Wu et al., 2023) make it trivial to wrap an LLM in an agent loop (plan → reason → act → observe).
@@
Although LLM-based agents have already been deployed with commerical usage in the real world, LLM-based agents remain subject to significant trustworthiness concerns and reliability challenges, due to LLM inherent issues such as errors (Yadkori et al., 2024), biases (Rettenberger et al., 2025), deceptions (Hubinger et al., 2024), hallucinations (Huang et al., 2025; Xu et al., 2025a), and misalignment (Shen et al., 2023), where the agent generates incorrect, not aligned with human value content or even harmful actions. While human-feedback reinforcement learning (RLHF) methods have been implemented to solve this alignment problem to bias reduction and harm minimization, the process provides no inherent guarantee of alignment with ethical values or desired behavioral norms (Barnhart et al., 2025). Researchers demonstrate “reverse alignment” attacks: with a few hours of fine-tuning, an open-weight model’s safety guardrails can be removed entirely (Yi et al., 2024). The proliferation of open models, while offering greater flexibility for customized model fine-tuning, intensifies these alignment challenges since anyone can fine-tune models in potentially harmful ways—even without awareness—operating outside the oversight of model companies and traditional governance like law (Mo et al., 2023; Hu et al., 2025).
@@ 2.3. Trustless Computational Substrates
DePIN combined with TEE computation like Phala Network [^3] https://phala.network/, a CPU/GPU TEE-based decentralized cloud computing platform, illustrates how TEEs and DePIN combine to support DeAgents. Phala decouples computation from the main chain and runs smart contracts inside secure enclaves on worker nodes. This design ensures that even when off-chain workers are operated by random volunteers, the code executes privately and correctly, with results returned to the blockchain. Phala markets itself as a ”trustless cloud computing” [^4] https://docs.phala.network/tech-specs/blockchain service for Web3 because it provides cloud-like computation without requiring users to trust server operators. More broadly, such TEE-enabled DePIN infrastructures offer permissionless and tamper-resistant compute power. A DeAgent deployed on a network like Phala can perform heavy AI computations or store data in a secure, decentralized manner, protected from both malicious hosts and external censors. The notion of trustless substrates thus encompasses both the ledger (blockchain) and the off-chain execution layer (TEE networks), creating a foundation where autonomous agents can operate free from centralized control or tampering.
@@ 2.5. Sociotechnical AI Governance
Recent work, including those shared at the Sociotechnical AI Governance session at CHI 2025, calls for sociotechnical approaches that integrate human and organizational context into AI governance (Feng et al., 2025b). Sociotechnical governance frameworks emphasize participatory oversight, transparency, and the inclusion of diverse stakeholders in decisions about AI agents. HCI scholars argue that social infrastructures (e.g. organizational culture, user values, regulatory norms) critically shape how AI systems should be governed (Dobbe and Wolters, 2024). Accordingly, researchers have begun promoting more democratic and community-driven governance processes. For example, Feng et al. (Feng et al., 2025b) also advocate moving beyond top-down control toward deliberative, participatory governance that empowers those impacted by an AI system to actively influence its design, evaluation, and policies. Ensuring trust and legitimacy also requires aligning AI behavior with human values and expectations. Bridging the “sociotechnical gap” in explainable AI is one noted challenge: there is often a divide between what technical explanations provide and the social context users need for those explanations to be meaningful (Ehsan et al., 2023).
The rise of DeAgents also surfaces critical challenges of governance: How do we steer or control autonomous agents that are deployed on open networks? By design, decentralized agents resist traditional governance. They are open-source (anyone can copy the code) and permissionlessly deployed (no central platform can easily shut them down), which can make them effectively ungovernable once released. This property echoes the governance dilemmas faced in other decentralized technologies (for example, BitTorrent or cryptocurrency protocols themselves), but with the added twist that these agents can make decisions and take actions in the world.
@@ 3.1.1. Semi-structured Interview
We conducted eight semi-structured interviews with people who had experiences deploying decentralized agents. Participants were recruited through outreach emails to known DeAgent developers and via snowball sampling based on referrals from initial interviewees. To protect participants’ privacy, we use generalized descriptions of their roles and affiliations rather than disclosing specific details. Basic background information about participants is provided in Table 1.
@@
P8 Researcher in Academia
@@ 3.1.2. Public Videos
We collected and transcribed eight public recordings from the Agentic Ethereum 2025 Summit [^7] Public link on YouTube: https://www.youtube.com/playlist?list=PLXzKMXK2aHh6DrYO4ORk8wzQ5wCZDs0mW, featuring talks by prominent community thought leaders. The recordings range from 17 to 56 minutes in length and include both open talks and panel discussions. These recordings were selected for two main reasons: 1) the summit not only addressed technical and marketing topics, but also emphasized governance opportunities and challenges about decentralized AI issues, which aligns with the aim of our research. 2) The speakers are key stakeholders with significant influence in the community, offering real-world insights and high-level perspectives on decentralized AI governance. Basic information of the speakers and video topic is provided in Table 2
@@
V1 Vitalik Buterin Co-founder of Ethereum Foundation The Promise & Challenge of Crypto + AI Applications
V2 Shaw Walters Co-founder, ElizaOS Why Should Ethereum Care About AI?
V3 Ed Felten Illia Polosukhin Co-founder Offchain Labs Co-founder of NEAR Protocol Where AI and Blockchain Intersect
V4 Tomasz Stanczak Founder of Nethermind ChaosChain and Agentic Governance
@@
V6 Erik Voorhees Founder of Venice Private AI
V7 Dmarz Jasper Zhang Davide Crapis Andrew Miller Dcbuilder Chase Allred Mark Tyneway Moderator at Flashbots Co-founder of Hyperbolic Co-founder of PIN AI Visiting Researcher at Flashbots Developer at World Engineer at Offchain Labs Developer at OP Labs Infra for Agentic Apps
V8 Kevin Owocki Tina Zhen Ken Ng Linda Xie Shaw Vitalik Buterin Founder of Gitcoin Contributor at Flashbots Developer at Uniswap Foundation Co-founder of Scalar Capital Co-founder of ElizaOS Co-founder of Ethereum Foundation Putting the Autonomy/Automation Back into DAOs
Table 2. Information of Videos in Agentic Ethereum 2025 Summit Playlist

### v2

# Trustless Autonomy: Understanding Motivations, Benefits and Governance Dilemmas in Self-Sovereign Decentralized AI Agents

@@ Abstract.
The recent trend of self-sovereign Decentralized AI Agents (DeAgents) combines Large Language Model (LLM)-based AI agents with decentralization technologies such as blockchain smart contracts and trusted execution environments (TEEs). These tamper-resistant trustless substrates allow agents to achieve self-sovereignty through ownership of cryptowallet private keys and control of digital assets and social media accounts. DeAgents eliminate centralized control and reduce human intervention, addressing key trust concerns inherent in centralized AI systems. This contributes to social computing by enabling new human cooperative paradigm ”intelligence as commons.” However, given ongoing challenges in LLM reliability such as hallucinations, this creates paradoxical tension between trustlessness and unreliable autonomy. This study addresses this empirical research gap through interviews with DeAgents stakeholders—experts, founders, and developers—to examine their motivations, benefits, and governance dilemmas. The findings will guide future DeAgents system and protocol design and inform discussions about governance in sociotechnical AI systems in the future agentic web.
@@ 1. Introduction
Deploying DeAgents on trustless secure enclaves, such as smart contracts and TEEs, inherently creates paradoxical tensions around governance risks due to diminished human oversight. Proponents envision that self-sovereign DeAgents can help moderate decentralized autonomous organizations (DAO) (Hinkes, 2024), manage financial portfolios (Ante, 2024), and etc., aiming to mitigate human error and centralized corruption (Trequattrini et al., 2024; Odilla, 2023). For example, ai16z is the first AI venture capital fund that is fully managed by DeAgent with recommendations from DAO members (ai1, 2024). However, while trustless infrastructure can guarantee tamper-resistant autonomous execution (Jauernig et al., 2020), the AI agent itself may not be inherently trustworthy (Liu et al., 2024). Large language models are prone to errors (Yadkori et al., 2024), biases (Rettenberger et al., 2025), deceptions (Hubinger et al., 2024), hypersensitive to nudges (Cherep et al., 2025), hallucinations (Huang et al., 2025; Xu et al., 2025a), and misalignment (Shen et al., 2023). The additional challenge of ”locking” AI agents in secure enclaves to prevent external tampering arguably amplifies potential harms if the agents’ decisions are flawed or hacked (Patlan et al., 2025)—especially without a kill switch (Seneviratne, 2024). This tension between trustlessness and autonomy fundamentally challenges the governability (Hu et al., 2025), accountability (Novelli et al., 2024) and responsibility (Dignum, 2019) of sociotechnical AI systems (Kudina and van de Poel, 2024; Dobbe and Wolters, 2024; Feng et al., 2025b).
@@
  • RQ2 (Emerging Social Dynamics): What roles do DeAgents come to play in human cooperation and social computing and how do people perceive them?
  • RQ3 (Challenges and Governance Implications): What challenges arise in the real-world deployment of DeAgents, and how do stakeholders navigate governance, oversight, and accountability within a trustless autonomous system?
@@
  • We presented empirical findings showing that the push for self-sovereign AI agents on trustless substrates is driven by sociotechnical / political factors like decentralization, privacy concerns, censorship resistance, and community ownership—even as AI reliability issues remain unresolved. Our study further identified DeAgents’ social impacts, including the evolving role they play, where the attention and trust they derived from. We summarized the difference between traditional AI agents and DeAgents.
  • We identified the paradoxical governance dilemma that arises when potentially unreliable AI agents operate within trustless infrastructure, revealing the complex tensions between trusting humans versus trusting AI in delegating agency to autonomous systems. Moreover, we found that the decentralized nature of DeAgents is the foundation to function as “commons” for social good, as it is manipulation-free.
  • We proposed design and governance implications that potentially address governance dilemma and emphasize thoughtful protocol design for agentic web of trust, with flexible fail-safes serving as essential bridges between trustlessness and genuine trustworthiness.
@@ 2.1. Terminology used in this paper
The capacity of an agent to perceive, decide, and act toward its goals without continuous human oversight, also known as ”self-governance.” However, higher powers may still exist that can assume control over the agent (Hui, 2024; Feng et al., 2025a).
@@
Trustless Agent
Agents that incorporate a trust layer enabling participants to discover, select, and engage with agents across organizational boundaries without requiring pre-existing trust (Rossi et al., 2025).
Decentralized Autonomous Organization (DAO)
DAO is an organization managed in whole or in part by decentralized computer programs, with voting and finances handled through a decentralized ledger technology like a blockchain.
@@ 2.2. LLM-based Agents
Recent advancements in Large Language Models have given rise to autonomous agents (Wang et al., 2024; Xi et al., 2023) capable of performing various real-world tasks (Xu et al., 2024), such as automatic social media marketing (Deshmukh et al., 2025), scientific discovery (Ren et al., 2025; Gridach et al., 2025) and financial trading (Ding et al., 2024). These LLM-based agents typically share a common architectural framework (Luo et al., 2025) that includes profile definition, memory mechanisms (Zhang et al., 2024a), planning capabilities (Huang et al., 2024; He et al., 2025), action execution via context protocol and external tools (Hou et al., 2025), inter-agent collaboration (Mou et al., 2024), and agent evolution (Dai et al., 2024). Open-source agentic frameworks such as LangChain and AutoGen (Wu et al., 2023) make it trivial to wrap an LLM in an agent loop (plan → reason → act → observe).
@@
Although LLM-based agents have already been deployed with commerical usage in the real world, LLM-based agents remain subject to significant trustworthiness concerns and reliability challenges, due to LLM inherent issues such as errors (Yadkori et al., 2024), biases (Rettenberger et al., 2025), hypersensitive to nudges (Cherep et al., 2025), deceptions (Hubinger et al., 2024), hallucinations (Huang et al., 2025; Xu et al., 2025a), and misalignment (Shen et al., 2023), where the agent generates incorrect, not aligned with human value content or even harmful actions. While human-feedback reinforcement learning (RLHF) methods have been implemented to solve this alignment problem to bias reduction and harm minimization, the process provides no inherent guarantee of alignment with ethical values or desired behavioral norms (Barnhart et al., 2025). Researchers demonstrate “reverse alignment” attacks: with a few hours of fine-tuning, an open-weight model’s safety guardrails can be removed entirely (Yi et al., 2024). The proliferation of open models, while offering greater flexibility for customized model fine-tuning, intensifies these alignment challenges since anyone can fine-tune models in potentially harmful ways—even without awareness—operating outside the oversight of model companies and traditional governance like law (Mo et al., 2023; Hu et al., 2025).
@@ 2.3. Trustless Computational Substrates
DePIN combined with TEE computation like Phala Network [^3] https://phala.network/, Oasis, a CPU/GPU TEE-based decentralized cloud computing platform, illustrates how TEEs and DePIN combine to support DeAgents. Phala decouples computation from the main chain and runs smart contracts inside secure enclaves on worker nodes. This design ensures that even when off-chain workers are operated by random volunteers, the code executes privately and correctly, with results returned to the blockchain. Phala markets itself as a ”trustless cloud computing” [^4] https://docs.phala.network/tech-specs/blockchain service for Web3 because it provides cloud-like computation without requiring users to trust server operators. More broadly, such TEE-enabled DePIN infrastructures offer permissionless and tamper-resistant compute power. A DeAgent deployed on a network like Phala can perform heavy AI computations or store data in a secure, decentralized manner, protected from both malicious hosts and external censors. The notion of trustless substrates thus encompasses both the ledger (blockchain) and the off-chain execution layer (TEE networks), creating a foundation where autonomous agents can operate free from centralized control or tampering.
@@ 2.4. The Emergence of DeAgents
2.5. Social Impact of Decentralization and Autonomous AI
2.5.1. Social Dynamics Change and Autonomous Agents
In the past decade, HCI/CSCW communities have investigated how autonomous agents reshape social dynamics. One stream of research treats agents as proxies to simulate human behavior for prototyping social computing systems or social experiments  (Park et al., 2022, 2024, 2023). Others focus on agents as service providers across sensitive or high-stakes domains. For instance, conversational agents have been developed to support mental health counseling (Qiu et al., 2025; Aleem et al., 2024) and educational purposes, acting as partners in collaborative learning or dialogue (Bendarkawi et al., 2025; Muñoz Cardona et al., 2021). Beyond domain-specific services, researchers also evaluate human–agent interactions in role-play settings to understand performance and social consequences (Höhn et al., 2024). Despite these contributions, most existing systems are developed and deployed within centralized infrastructures. The decision models, data pipelines, and design priorities of such agents are therefore largely determined by the values of the central entities that control them. This centralization raises questions about how power, agency, and value alignment are negotiated in human–agent social dynamics.
2.5.2. Social Dynamics Change and Decentralized Technologies
In addition, HCI and CSCW researchers have examined how decentralization alters social dynamics by redistributing authority and enabling alternative forms of interaction. Decentralized infrastructures delegate decision-making and action away from a central authority, foregrounding value-driven design practices that promote social transformation (Walker, 2020). One prominent line of work explores decentralized autonomous organizations (DAOs) as collective frameworks for democratic decision-making, applied to contexts ranging from community planning to co-design processes (Yokoyama et al., 2025; Xiang et al., 2023). Decentralization has also been leveraged to reconfigure trust. For example, decentralized identity management offers new mechanisms for establishing trust among individuals, groups, and technologies (Sharma et al., 2024; Rankin et al., 2020). Financial transactions and payment infrastructures have been reconceptualized through decentralized systems, producing new paradigms for exchange (Nissen et al., 2018; Helleiner, 2000) and tamper-proof records for notary or legal processes (Elsden et al., 2018). In the social media domain, decentralized infrastructures are investigated as a means of balancing power asymmetries, granting user agency through decentralized content moderation (Zhang et al., 2024b, 2025; Hwang et al., 2025) and feed curation (Liu et al., 2025; Feng et al., 2024). Likewise, decentralized communication platforms such as Matrix offer stronger privacy protections through federated and encrypted messaging. Most of this prior work conceptualizes decentralized technologies as facilitators of community or group interaction. More recent research, however, begins to envision autonomous agents as independent entities within decentralized systems. In this work, we extend these trajectories by investigating the social impact of agents that are both autonomous and deployed on decentralized infrastructures, and the novel dynamics that arise from this intersection.
@@ 2.6. Sociotechnical AI Governance
Recent work, including those shared at the Sociotechnical AI Governance session at CHI 2025, calls for sociotechnical approaches that integrate human and organizational context into AI governance (Feng et al., 2025c). Sociotechnical governance frameworks emphasize participatory oversight, transparency, and the inclusion of diverse stakeholders in decisions about AI agents. HCI scholars argue that social infrastructures (e.g. organizational culture, user values, regulatory norms) critically shape how AI systems should be governed (Dobbe and Wolters, 2024). Accordingly, researchers have begun promoting more democratic and community-driven governance processes. For example, Feng et al. (Feng et al., 2025c) also advocate moving beyond top-down control toward deliberative, participatory governance that empowers those impacted by an AI system to actively influence its design, evaluation, and policies. Ensuring trust and legitimacy also requires aligning AI behavior with human values and expectations. Bridging the “sociotechnical gap” in explainable AI is one noted challenge: there is often a divide between what technical explanations provide and the social context users need for those explanations to be meaningful (Ehsan et al., 2023).
The rise of DeAgents also surfaces critical challenges of governance: How do we steer or control autonomous agents that are deployed on decentralized permissionless networks? By design, decentralized agents resist traditional governance. They are open-source (anyone can copy the code) and permissionlessly deployed (no central platform can easily shut them down), which can make them effectively ungovernable once released. This property echoes the governance dilemmas faced in other decentralized technologies (for example, BitTorrent or cryptocurrency protocols themselves), but with the added twist that these agents can make decisions and take actions in the world.
@@ 3.1.1. Semi-structured Interview
We conducted thirteen semi-structured interviews with people who had experiences deploying decentralized agents. Participants were recruited through outreach emails to known DeAgent developers and via snowball sampling based on referrals from initial interviewees. To protect participants’ privacy, we use generalized descriptions of their roles and affiliations rather than disclosing specific details. Basic background information about participants is provided in Table 1.
@@
P8 Decentralized AI Researcher
P9 Developer at a Decentralized Agentic Cloud Company
P10 Co-founder of a Decentralized Agentic Cloud Company
P11 Decentralized Protocol Researcher
P12 Builder of a Decentralized Protocol
P13 Co-founder of a Machine Economy Investment Company
@@ 3.1.2. Public Videos
We collected and transcribed ten public recordings from the Agentic Ethereum 2025 Summit (V1 - V8) [^7] Public link on YouTube: https://www.youtube.com/playlist?list=PLXzKMXK2aHh6DrYO4ORk8wzQ5wCZDs0mW and the recommendation of our study participants (V9 and V10), featuring talks by prominent community thought leaders. The recordings range from 17 to 56 minutes in length and include both open talks and panel discussions. These recordings were selected for two main reasons: 1) the summit not only addressed technical and marketing topics, but also emphasized governance opportunities and challenges about decentralized AI issues, which aligns with the aim of our research. 2) The speakers are key stakeholders with significant influence in the community, offering real-world insights and high-level perspectives on decentralized AI governance. Basic information of the speakers and video topic is provided in Table 2
Table 2. Information of Public Videos Analyzed in the study
@@
V1 Vitalik Buterin Co-founder of Ethereum Foundation
The Promise and Challenge of
Crypto + AI Applications
V2 Shaw Walters Co-founder of ElizaOS
Why Should Ethereum Care About AI?
V3
Ed Felten
Illia Polosukhin
Co-founder Offchain Labs
Co-founder of NEAR Protocol
Where AI and Blockchain Intersect
V4 Tomasz Stanczak Founder of Nethermind
ChaosChain and Agentic Governance
@@
V6 Erik Voorhees Founder of Venice Panel: Private AI
V7
Dmarz
Jasper Zhang
Davide Crapis
Andrew Miller
Dcbuilder
Chase Allred
Mark Tyneway
Moderator at Flashbots
Co-founder of Hyperbolic
Co-founder of PIN AI
Visiting Researcher at Flashbots
Developer at World
Engineer at Offchain Labs
Developer at OP Labs
Panel: Infra for Agentic Apps
V8
Kevin Owocki
Tina Zhen
Ken Ng
Linda Xie
Shaw Walters
Vitalik Buterin
Co-founder of Gitcoin
Co-founder at Flashbots
Developer at Uniswap Foundation
Co-founder of Scalar Capital
Co-founder of ElizaOS
Co-founder of Ethereum Foundation
Panel: Putting the Autonomy/
Automation Back into DAOs
V9 Davide Crapis Marco De Rossi Shaw Walters Co-creator of ERC-8004 Co-creator of ERC-8004 Co-founder of ElizaOS Trustless Agents — ERC-8004 Deep Dive
V10
David Dao
Sejal Rekhan
Co-founder of DeepGov
DeepGov: Configurable AI Politicians
for Capital Allocation and Governance